OpenStudy (anonymous):

Find the point on the sphere x^2 +y^2 + z^2 = 9 that is closest to the point (2,3,4) (presumably using lagrange multipliers, since thats what the chapter entails lol)

OpenStudy (blockcolder):

You have to minimize the distance between a point on the sphere and the point (2,3,4), with the constraint that the point lies on the sphere.

OpenStudy (anonymous):

how do i do that? :(

OpenStudy (dumbcow):

distance function is: \[\sqrt{(x-a)^{2} +(y-b)^{2} +(z-c)^{2}}\] let a,b,c be point (2,3,4) and x,y,z point on sphere \[z = \sqrt{9-x^{2}-y^{2}}\] \[d = \sqrt{(x-2)^{2}+(y-3)^{2}+(\sqrt{9-x^{2}-y^{2}}-4)^{2}}\] set partial derivatives equal to 0 and solve for x,y

OpenStudy (anonymous):

where did you get the last 4 from? and is there a way to also do this with lagrange?

OpenStudy (dumbcow):

from the point (2,3,4) and probably...the Lagrange multipliers may come into play later in the problem, i've forgotten how to use those...something with matrices

OpenStudy (anonymous):

so.. how do i set partials when everything is under a root? :\ and the lagrange multiplier is just: f(x,y,z) - lambda (g(xyz) - k)

OpenStudy (dumbcow):

you just take the derivative...its a lot of the chain rule :)

OpenStudy (anonymous):

would you mind showing me the steps for just one of them, like Fx? the only thing i can think of is \[\sqrt{2(x-2)}\]

OpenStudy (dumbcow):

do you know the solution? i want to make sure i did it right first...i didn't use langrange multiplier at all so this won't help you for your class too much :(

OpenStudy (anonymous):

8/3, according to the book

OpenStudy (dumbcow):

? it should be a point

OpenStudy (anonymous):

just kidding, wrong answer ^.^ the correct one is \[6/\sqrt{29}, 9/\sqrt{29}, 12/\sqrt{29}\]

OpenStudy (dumbcow):

ok i got the right answer but doing it a different way

OpenStudy (dumbcow):

here you go..haha http://en.wikipedia.org/wiki/Lagrange_multiplier

OpenStudy (anonymous):

yeahhh i get how to use them, just not for this problem :(

OpenStudy (dumbcow):

hmm try \[f(x,y,z) = \sqrt{(x-2)^{2} +(y-3)^{2}+(z-4)^{2}}\] \[g(x,y,z)-c = x^{2}+y^{2}+z^{2}-9\] then \[\rightarrow \sqrt{(x-2)^{2} +(y-3)^{2}+(z-4)^{2}} +\lambda( x^{2}+y^{2}+z^{2}-9)\]

OpenStudy (dumbcow):

now you have to take 4 partial derivatives..fun

OpenStudy (anonymous):

ohhhhhh okay :o but.. i still don't understand how to even begin going about getting a partial derivative while everything is under the square root :( :(

OpenStudy (dumbcow):

\[f(x) = \sqrt{u(x)}\] \[f'(x) = \frac{1}{2\sqrt{u(x)}}*u'(x)\] chain rule

OpenStudy (dumbcow):

i'll do Fx the derivative inside root is just 2(x-2) right multiply that by derivative of sqrt(u) ...where u is that whole big function \[F_x = \frac{2(x-2)}{2\sqrt{u}} +2 \lambda x\]

OpenStudy (anonymous):

i'm kind of having a hard time understanding what you're doing, but let me try..is this correct? \[f_y=2(y-3)/2\sqrt{u}\]

OpenStudy (anonymous):

+2lambda y

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

so then \[f_z = 2(z-4)/2\sqrt{u} + 2z \lambda \]

OpenStudy (dumbcow):

do you know why the sqrt(u) is on bottom? \[d/dx \sqrt{x} = x^{1/2} = \frac{1}{2}x^{-1/2} = \frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

OHHHHHHHHHH that totally makes so much more sense now LOL

OpenStudy (dumbcow):

good...also you can simply all of those by cancelling out the 2's

OpenStudy (dumbcow):

not in lambda part though

OpenStudy (anonymous):

wait another question ... so like lets pretend that i did the chain rule but it WASN'T 2(x-2)/2(u) but it was 2(x-2)/(u) right? and then i had to solve for x... i know you're not supposed to touch the inside when doing the chain rule, but when solving for x, after doing the chain rule.. would i be allowed to distribute then?

OpenStudy (dumbcow):

yes after differentiating you can distribute if you need to

OpenStudy (anonymous):

ohh okay! great! thank you sooooo much!!!! i wish i could like.. give you two medals, or fan you again lol

OpenStudy (dumbcow):

haha no problem...good luck , you may want to review derivative rules before the test

OpenStudy (anonymous):

so the denominator is 2sqrt {(x-2)^2+y-3^2+(z-4)^2, right? do i need to write that out for fx fy and fz? and what can i do with it when i'm solving for stuff? :o

OpenStudy (dumbcow):

its up to you...since its same for each one you could leave it as "u" or some other variable until you have to break it up i think for this problem it just ends up cancelling out so it doesn't matter

OpenStudy (anonymous):

wait.. is it supposed to be f(x) - lambda g(x) or f(x) + lambda g(x)?

OpenStudy (dumbcow):

it says "+" do you use "-" for minimization problems?

OpenStudy (dumbcow):

hold on let me solve it and i'll find out

OpenStudy (anonymous):

i actually have no idea, i just know the formula i've been using is -lambda :o but that's been for finding local max and min, not sure if that counts as minimization =\ and they don't really explain how to do this type of problem in the book -_-"

OpenStudy (anonymous):

btw, is this right? \[2x \lambda = (x-2)/\sqrt{u} , x = (x-2)/ 2 \lambda \sqrt{u}\]

OpenStudy (dumbcow):

ok using "+" works, i get same answers as before looks right...change the sign though

OpenStudy (anonymous):

okay! thank you!! :D

OpenStudy (dumbcow):

your welcome

OpenStudy (anonymous):

so... i'm kind of stuck again :x i got x = -x+2 / 2λ√u y = -y+3 / 2λ√u z = - -z+4/2λ√u and am not sure how to proceed =\

OpenStudy (anonymous):

sorry :(

OpenStudy (dumbcow):

haha no problem...well you need to isolate x, right now you have an x on both sides 2 ways to do this: 1. split right side into 2 fractions, then move x term over 2. multiply the denominator back over to left side, then add the x over...factor out the x..

OpenStudy (anonymous):

i just tried to do it in my head real quick, and do they both yield the same answer? :o

OpenStudy (dumbcow):

yes they do...no matter how you do it x can only equal one expression right

OpenStudy (anonymous):

haha oh right... -_-" i knew that..

OpenStudy (anonymous):

okay so this may be a silly question but ... lets call 2λ√u m? i split it into two to get x = -x/m + 2/m, moved the x over to get x + x/m = 2/m so then i factor it? to get x (1+1/m) = 2/m ? now.. if i divide by 1+1/m .... ... i'm not sure how to do this :X

OpenStudy (dumbcow):

ahh yes, lot of algebra involved here you need to combine 1+ 1/m into 1 fraction, then to divide, multiply by its reciprocal

OpenStudy (anonymous):

then.. would it be... 1m/m + 1/m? o.O

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

would that make it 2m/m? would the m's cancel out? to just be 2?! :o

OpenStudy (dumbcow):

uh oh , you are missing something...but yes you will have the m's cancel m/m + 1/m = (m+1)/m reciprocal -> m/(m+1)

OpenStudy (anonymous):

OH lol -_-" that was silly haha... so then is x = 2/m+1?

OpenStudy (dumbcow):

sorry..yes thats correct

OpenStudy (anonymous):

yes! okay :D great! you're my hero. x = 2/m+1 , y = 3/m+1, and z = 4/m+1?!

OpenStudy (dumbcow):

good...now plug them into last equation which is just g(x,y,z)-c and solve for lambda

OpenStudy (anonymous):

so i immediately factored out m to get m(-2(2) - 2(3) - 2(4) which is m(-18) = 9 -18 = 9/m, divide both sides by 9 i get -2 = 1/m, or -2 = 1/2λ√u+1 i'm not sure what else i can do to try to get lambda alone :(

OpenStudy (dumbcow):

what equation are you using? it should be partial wrt lambda --> x^2 +y^2 +z^2 -9 = 0

OpenStudy (dumbcow):

and be careful...denominator is 1+m so you can't factor out a m

OpenStudy (dumbcow):

oh nevermind, i see what you did....m is "m+1" :)

OpenStudy (anonymous):

wait i'm using g(x)?? not g'x?!? O.O i thought i had to find fx, fy, fz, and then f lamdba =\

OpenStudy (anonymous):

oh.. i see what i did wrong -_-"

OpenStudy (dumbcow):

think , what is f lambda ?

OpenStudy (anonymous):

yeah sorry, m = 2λ√u and we'll make n = 2λ√u+1 just so it makes more sense haha.. so plugging everything in i get 4/n + 9/n + 16/n = -9, can i multiply both sides by n?

OpenStudy (dumbcow):

yes...the -9 should be +9 when you move it over also the denominator is squared...so technically it should be n^2

OpenStudy (anonymous):

oh! i totally forgot about the square -_-" so then is it 4 lambda u? which will make it 28 = 36 lambda^2 u?

OpenStudy (dumbcow):

lets see...you forgot about the +1 ....n = 1+2lambda*sqrt(u) 28 should be 29....4+9+16

OpenStudy (anonymous):

i subtracted 1 from both sides, am i not allowed to do that? :(

OpenStudy (anonymous):

btw.. does u ever cancel? :o

OpenStudy (dumbcow):

29 = 9n^2 take square root sqrt29 = 3n sqrt29/3 = n = 1+2lambda*sqrt(u)

OpenStudy (dumbcow):

then subtract 1

OpenStudy (dumbcow):

and yes the u will cancel when we plug lambda back in to solve for x,y,z

OpenStudy (anonymous):

ohhhhhh okay, order of operations, lol that's right. so then is it: \[-1+\sqrt{29}/6\sqrt{u} = \lambda\]

OpenStudy (dumbcow):

is that (-1 +sqrt29) ? should be (-3+sqrt29) when you combine fractions...change 1 to 3/3

OpenStudy (anonymous):

OH i totally did that at first andthen was like nahh... lol -_-" okay -3+! is the denominator correct? :o

OpenStudy (dumbcow):

haha yes you got it

OpenStudy (anonymous):

YES!!!!! :D :D : D :D omg thank you soooooooooooo much!!!!!!!!! hey, btw, are you a girl or guy? lol

OpenStudy (anonymous):

i just realized that sounds creepy lol. i want to write you a testimony and would like to get the sex correct haha

OpenStudy (dumbcow):

not done yet...still have to plug in lambda to get x,y,z and im a guy...guess i could add that to my profile

OpenStudy (anonymous):

haha with a name like dumbcow that's actually pretty hard to guess. it's not fair that plugging lambda in makes things even more complicated

OpenStudy (dumbcow):

its not too bad...the sqrt(u) will cancel , just be careful with combining fractions and simplifying actually i will say this method turns out to be less messy than my original way

OpenStudy (anonymous):

actually, jk, maybe not... is it 1/root(29)? :o

OpenStudy (dumbcow):

close...haha you have the solutions remember

OpenStudy (anonymous):

yeah but i refuse to look at them :o and i'm not plugging into the original equation yet, i'm just trying to solve for x y and z, right?

OpenStudy (dumbcow):

right x = 2/(1+2lambda*sqrt(u)) now take the expression for lambda and plug it in

OpenStudy (anonymous):

so, i must confess, i'm not exactly sure how to properly plug it in haha but this is what i did: 2/2 x (-3+root(29)/6rootu) x root(u)+1 = 2 / -6 +2root(29)6root(u)) x root(u) +1 , i think the 6 and root u cancel out to get 2 / -1+2root(29)+1, the 1's cancel out to get 2/2root(29) and the 2's reduce to get 1/root(29) :o

OpenStudy (dumbcow):

ok you're mistake is the 6's do not cancel because the whole numerator is (-6+2root29) if it was (-6+6root29) then you could factor out a 6, then 6's cancel in this case, however, the most you can factor out is a 2 --> 2(-3+root29)/6 = (-3+root29)/3

OpenStudy (anonymous):

oh...so then it's... but wait.. this is already a denominator.. so what happens with a denominator of a denominator?! i have 2 / (-2+root(29)/3) is that right?

OpenStudy (dumbcow):

careful with those fractions \[\frac{\sqrt{29}-3}{3} +1 \neq \frac{\sqrt{29}-2}{3}\]

OpenStudy (anonymous):

dammit. 2/ (3+root(29)/6) ?!

OpenStudy (dumbcow):

Note: when you have to divide by a fraction...multiply by its reciprocal

OpenStudy (dumbcow):

no thats not it...its alright just think get common denominator right change the 1 to 3/3

OpenStudy (anonymous):

oh lol wait i had it over 6 for some reason...its 0 ^.^ 2 / (root(29)/6) you said to multiply by the reciprocal? so then should it be 2/6root(29)?

OpenStudy (dumbcow):

haha its still over 6 :) reciprocal means flip the fraction....root(29)/3 -> 3/root(29)

OpenStudy (anonymous):

why.. why is that six.. still there -_-"..... i'm sorry, you must think i'm sooo stupid lol. please don't hate me... wait so .. am i multplying the entire thing by the reciprocal? so like 2/(root29/3) x 3/root29?

OpenStudy (anonymous):

or am i mutiplying 2 /root(29)/3 by 3/root(29)/2? lol